If x=a(cos t+cos 2t) & y=b(sin t+sin 2t) then find an equation without t.

Grade:11

1 Answers

Arun
5 years ago
Add b²x² and a²y²
 
Hence
b² x²+a² y²  = a² b² + a²b² + 2a²b² cos t = 2a² b² (cos t +1)
 
Now
 
If x/y is find out
 
Then 
 
x/y = a(2cost -1) ( cost +1) / b(sint ) (2 cos t +1)
 
Here you can find out cos t +1
 
And then you can find the relation between x and y